OpenStudy (anonymous):

....!Help Please!.... Solve the system of equations by substitution. What is the solution for x? *-> How do I solve this? O.o 2x + y = 1 4x + 2y = -1 x = 0 x = 2 There is no x value as there is no solution. x can be any value, as there is an infinite number of solutions.

OpenStudy (eric_d):

rearrange them 2x + y = 1 y=1-2x............... 4x + 2y = -1 2y=-1-4x y=-1/2-2x................. 1-2x=-1/2-2x -2x+2x=-1/2-1 0= -1.5 There is no x value.....

OpenStudy (anonymous):

Thank You so much! @eric_d c:

OpenStudy (eric_d):

Np :)